# Robust nonlinear state-feedback control of second-order systems

### v2

@@ III-B Uniform asymptotic stability
The origin for the system (3) is uniformly asymptotically stable for any $x(0)\in\mathbb{X}$ and $\|d\|_{\infty}\leq D$ if
$\gamma>D^{1.5}+D+\frac{1}{2}.$ (6)
Proof:
As in the Appendix for the system (9), consider for (3) a strict Lyapunov function candidate
@@
with $z_{1}=\sqrt{|x_{1}|}\,\mathrm{sign}(x_{1})$, which is positive definite for $\sqrt{2\gamma}>\varepsilon$. This Lyapunov function is continuously differentiable for $x\in\mathbb{X}\setminus\{0\}$. Taking the time derivative of $V$ on this set one obtains
@@
both expressions in the brackets on the right hand-side of $\dot{V}$ are positive under (6). Hence, $\dot{V}<0$ for $V\neq 0$ in $\mathbb{X}$ that completes the proof. ∎
The low bound of the control gain $\gamma$, which is satisfying the parametric condition (6), is shown as a function of the maximal disturbance magnitude, i.e., $D$, in Fig. 3.
@@
III-C Finite-time convergence
In order to prove the finite-time convergence (see, e.g., [4] for details) of the system (3), which can be supposed from the numerical results given in Fig. 2 (a), we make further use of the strict Lyapunov function $V$ given in (7).
@@ III-C Finite-time convergence
Note that $(d\,\text{sign}(x_{1})-\gamma)|x_{1}|^{-1}<0$ due to (6), which implies that for any $x(0)\in\mathbb{X}$ and $\|d\|_{\infty}\leq D$ there is $t_{1}\geq 0$ such that $\zeta(t)<0$ for all $t\geq t_{1}$, i.e., the trajectories enter in the second or the fourth quadrants after finite transients, where the variable $\zeta$ stays separated with zero. Next, consider another auxiliary variable $z(t)=x_{2}^{2}(t)|x_{1}(t)|^{-1}$, which admits the dynamics:
@@
For $t\geq t_{1}$ and taking into account (6), $\text{sign}(x_{2})(d-\gamma\text{sign}(x_{1}))>0$ out the origin (the lines $x_{1}=0$ and $x_{2}=0$ can be reached only there). Hence, for any $x(0)\in\mathbb{X}$ and $\|d\|_{\infty}\leq D$, there is $t_{2}\geq 0$ such that $z(t)\geq\theta>0$ for all $t\geq t_{2}$ (while the trajectories do not converge to zero), where $\theta$ depends only on the choice of $\gamma$. So, there are two possibilities, either the trajectories approach the origin in a finite time, or the variable $z(t)$ becomes strictly positive and separated with zero, which returning to the Lyapunov function $V(x)$ from (7) and the estimate (8) implies for $t\geq t_{2}$:
@@
for some $\kappa>0$. The latter estimate implies a finite-time convergence of $V$ to zero, according to [4]. ∎
III-D Simulations
The controlled output behavior of (3) is exemplary visualized in Fig. 4 (a). Here a periodic disturbance $d(t)=50\sin(20\pi t)$ is used while the control gain is assigned as $\gamma=100$. Both the disturbance signal and the control value are shown opposite to each other in Fig. 4 (b).
Figure 4: Output response of the perturbed control system (3) in (a), and the control signal versus disturbance value in (b).
@@ IV Experimental control results
The proposed control (3) is experimentally evaluated on the electro-magneto-mechanical actuator system with one translational degree of freedom, see Fig. 5. The system dynamics can be reduced to the model (1).
@@
The step response of the control (3) (denoted further as NLG due to the nonlinear feedback gains) is compared once with a standard proportional-derivative (PD) control and once with the sub-optimal nonlinear damping control [14] (denoted further as SOND).
@@
param. \  cont. $\mathrm{PD}_{1}$ $\mathrm{PD}_{2}$ $\mathrm{SOND}$ $\mathrm{NLG}$
@@
For a fair comparison, the SOND control is assigned with the same gaining factor $\gamma=750$ as the faster PD-control $\mathrm{PD}_{2}$. The feedback gain of the NLG-control $\gamma=25$ is intentionally set an order of magnitude lower than the feedback gains of the PD and SOND controllers. This is for demonstrating that the NLG-control can behave less aggressive during the transient response and, at the same time, achieve a much higher accuracy at steady-state. All experimentally measured control responses, to the same reference step of $y_{\mathrm{ref}}=8$ mm, are shown opposite each other in Fig. 6.
@@
From the response of the NLG-control, one can also recognize a nearly uniform behavior of $x_{2}(t)$ during the whole transient phase, cf. Fig. 2 (a). Note that the corresponding inclination of the $x_{1}(t)$ slope depends on the assigned $\gamma$ parameter. It is also worth noting that a residual non-zero, however minor, steady-state error of the NLG-control is due to the sensor’s bias and noise, while the latter is additionally impacting the used $x_{2}(t)$ signal, which is obtained via differentiation and filtering of the measured $x_{1}(t)$. The experimentally obtained signal $x_{2}(t)$ (when using the NLG controller) is also shown in Fig. 7 for the sake of clarity.
Figure 7: The velocity state $x_{2}(t)$ of NLG control experiment.
@@ V Summary
A novel robust nonlinear state-feedback control for second-order systems was introduced. The control law combines the ideas of the state-dependent nonlinear damping introduced in [14] and the sliding mode control theory, which equips the overall feedback law with an output-sign-dependent term able to compensate for bounded perturbations. The resulted control has only one tuning parameter, the overall output feedback gain, whose selection should be based on the upper bound of the perturbation magnitude, and this closed analytic dependence was shown in the paper. We provided a detailed analysis of uniform asymptotic stability property of the control system and finite-time convergence based on the proposed strict Lyapunov function. The control signal appears to be continuous during the transients, apart from the convergence instant, where the state value reaches the origin (similarly to some high order sliding mode control algorithms). Remarkable features of the proposed control are that: (i) it is equivalent (from the structural viewpoint) to a standard linear PD feedback control and, at the same time, (ii) it can guarantee (without additional integral action) for full compensation of the bounded matched disturbances in the second-order systems without overshoot, while (iii) providing a global finite-time convergence. The proposed control was already evaluated experimentally on an electro-magneto-mechanical actuator system with considerable bounded disturbances. Analysis of its robustness abilities against process and measurement noise and development of the methods for its discretization can be considered as directions of future research.

### v3

@@ III-B Analysis of stability
The origin for the system (3) is uniformly attractive for any $x(0)\in\mathbb{X}$ and $\|d\|_{\infty}\leq D$ if
$\gamma>D^{1.5}+D+\frac{1}{2},$ (5)
and the solutions are bounded for all $t\geq 0$.
Proof.
As in the Appendix for the system (.1), consider for (3) a strict Lyapunov function candidate
@@
with $z_{1}=\sqrt{|x_{1}|}\,\mathrm{sign}(x_{1})$, which is positive definite for $\sqrt{2\gamma}>\varepsilon>0$. This Lyapunov function is continuously differentiable for $x\in\mathbb{X}\setminus\{0\}$. Taking the time derivative of $V$ on this set one obtains
@@
both expressions in the brackets on the right hand-side of $\dot{V}$ are positive under (5). Hence, $\dot{V}(x,d)<0$ for $x\in\mathbb{X}\setminus\{0\}$ and $|d|\leq D$ that completes the proof. ∎
Remark 3.
A consequence of the result given in Theorem 2, which establishes forward invariance of the set $\mathbb{X}$ and global convergence of all solutions in (5) to the origin, is that it implies the absence of overshooting: the trajectories cannot cross the line $x_{1}=0$, and they reach it only at the origin.
The low bound of the control gain $\gamma$, which is satisfying the parametric condition (5), is shown as a function of the maximal disturbance magnitude, i.e., $D$, in Figure 3.
@@
III-C Finite-time convergence and control boundedness
In order to prove the finite-time convergence (see, e.g., [4] for details) of the system (3), which can be supposed from the numerical results given in Figure 2 (a), we make further use of the strict Lyapunov function $V$ given in (6).
@@ III-C Finite-time convergence and control boundedness
Note that $(d\,\text{sign}(x_{1})-\gamma)|x_{1}|^{-1}<0$ due to (5), which implies that for any $x(0)\in\mathbb{X}$ and $\|d\|_{\infty}\leq D$ there is $t_{1}\geq 0$ such that $\zeta(t)<0$ for all $t>t_{1}$, i.e., the trajectories enter in the second or the fourth quadrants after finite transients, where the variable $\zeta$ stays separated with zero. Next, consider another auxiliary variable $z(t)=x_{2}^{2}(t)|x_{1}(t)|^{-1}$, which admits the dynamics:
@@
For $t>t_{1}$ and taking into account (5), $\text{sign}(x_{2})(d-\gamma\text{sign}(x_{1}))>0$ out the origin (the lines $x_{1}=0$ and $x_{2}=0$ can be reached only there). Hence, for any $x(0)\in\mathbb{X}$ and $\|d\|_{\infty}\leq D$, there is $t_{2}\geq 0$ such that $z(t)\geq\theta>0$ for all $t\geq t_{2}$ (while the trajectories do not converge to zero), where $\theta$ depends only on the choice of $\gamma$ and $D$. So, there are two possibilities, either the trajectories approach the origin in a finite time, or the variable $z(t)$ becomes strictly positive and separated with zero, which returning to the Lyapunov function $V(x)$ from (6) and the estimate (7) implies for $t\geq t_{2}$:
@@
for some $\kappa>0$. The latter differential inequality ensures a finite-time convergence of $V$ to zero according to [4]. ∎
Slightly extending the arguments used in the last proof we can substantiate the boundedness of the control (2):
Proposition 5.
In the conditions of Theorem 4, the amplitude of control (2) admits an upper estimate
$|u(t)|\leq\gamma+\max\{x_{2}^{2}(0)|x_{1}(0)|^{-1},2(D+\gamma)\},\;\forall t\geq 0.$
Proof.
Following the proof of Theorem 4 and considering an auxiliary Lyapunov function $\mathcal{V}(z)=0.5z^{2}$ we get the estimate on its time derivative for $t\geq 0$ while $z(t)\neq 0$ (until the origin is not reached, since it is the only invariant solution on the set $z=0$):
$\dot{\mathcal{V}}=\frac{|x_{2}|}{|x_{1}|}\,\Bigl{(}2\text{sign}(x_{2})\bigl{(}d-\gamma\,\text{sign}(x_{1})\bigr{)}z$
$-\bigl{(}2+\text{sign}(x_{1})\text{sign}(x_{2})\bigr{)}z^{2}\Bigr{)}$
$\leq\frac{|x_{2}|}{|x_{1}|}\,\Bigl{(}2(D+\gamma)|z|-z^{2}\Bigr{)}\leq\frac{|x_{2}|}{|x_{1}|}\,\Bigl{(}2(D+\gamma)^{2}-\mathcal{V}\Bigr{)},$
which implies the following boundedness estimate in the time domain: $\mathcal{V}(z(t))\leq\max\{\mathcal{V}(z(0)),2(D+\gamma)^{2}\}$ (the multiplier $\frac{|x_{2}|}{|x_{1}|}$ may be zero at the isolated time instant $t_{1}$ when $x_{2}(t_{1})=0$ only), or
$z(t)$ $\leq\max\{z(0),2(D+\gamma)\}$
for all $t\geq 0$. Note that $|u(t)|\leq\gamma+z(t)$ for all $t\geq 0$ in (2) that gives the required upper bound for the control. ∎
Moreover, since the set $z\leq 2(D+\gamma)$ is attractive and forward invariant (it contains the origin), it implies that for any trajectory approaching the origin (even initially $x_{1}(0)$ was arbitrary small) the control has an upper bound
$|u|\leq 2D+3\gamma,$
and the term $\frac{|x_{2}|x_{2}}{|x_{1}|}$ does not impose a singularity in the system (3). In addition, for the set of initial conditions with $x_{1}(0)\neq 0$ and $x_{2}(0)=0$ (i.e., when the control is applied to the system (1) in a steady state for $u(t)=d(t)=0$, $t\leq 0$), the amplitude of (2) is bounded by $2D+3\gamma$ for all $t\geq 0$.
Remark 6.
The proven asymptotic boundedness of $z$ and also its separation with zero, i.e., the forward attractiveness of the set $\theta\leq z\leq 2(D+\gamma)$ while the trajectory stays in $\mathbb{X}\setminus\{0\}$, implies that (3) admits an invariant cone in the second and third quadrants.
III-D Numerical simulations
Figure 4: Output response of the perturbed control system (3) in (a), and the control signal versus disturbance value by using: thresholding scheme (8) for zero in (b), and regularization scheme (9) in (c).
The controlled output behavior of (3) is exemplary visualized in Figure 4 (a). Here a periodic disturbance $d(t)=50\sin(20\pi t)$ is used while the control gain is assigned as $\gamma=100$. The most simple first-order forward Euler method (solver) is used. For numerical solvability of the closed-loop system dynamics (3), two implementation schemes are used. The first one
$u(t)=\begin{cases}-|x_{1}(t)|^{-1}\bigl{(}\gamma x_{1}(t)+|x_{2}(t)|x_{2}(t)\bigr{)}&|x_{1}(t)|\geq\mu\\ -\gamma\,\text{sign}\bigl{(}x_{1}(t)\bigr{)}&|x_{1}(t)|<\mu\end{cases},$ (8)
incorporates a threshold value $\mu>0$ for realizing numerical zero. The corresponding control signal is shown opposite the disturbance value in Figure 4 (b). For the sake of clarity, the same control signal but low-pass (LP) filtered with 3000 Hz cut-off frequency is also depicted. The second implementation scheme uses a regularization factor $\mu>0$, cf. [17], so that
$u(t)=-\Bigl{(}\bigl{|}x_{1}(t)\bigr{|}+\mu\Bigr{)}^{-1}\Bigl{(}\gamma x_{1}(t)+\bigl{|}x_{2}(t)\bigr{|}x_{2}(t)\Bigr{)}.$ (9)
The corresponding control signal is shown opposite the disturbance value in Figure 4 (c). Note that for both implementation schemes, a low numerical value $\mu=10^{-9}$ is assumed.
@@ IV Experimental control results
The proposed control (3) is experimentally evaluated on the electro-magneto-mechanical actuator system with one translational degree of freedom, see Figure 5. The system dynamics can be reduced to the model (1).
@@
The step response of the control (3) (denoted further as NLG owing to the nonlinear feedback gains) is compared once with a standard proportional-derivative (PD) control and once with the sub-optimal nonlinear damping control [15] (denoted further as SOND).
@@
parameter \  control $\mathrm{PD}_{1}$ $\mathrm{PD}_{2}$ $\mathrm{SOND}$ $\mathrm{NLG}$
@@
Note that for a fair comparison, the SOND control is assigned with the same gaining factor $\gamma=750$ as the faster PD-control $\mathrm{PD}_{2}$. The feedback gain of the NLG-control $\gamma=25$ is intentionally set an order of magnitude lower than the feedback gains of the PD and SOND controllers. This is for demonstrating that the NLG-control can behave less aggressive during the transient response and, at the same time, achieve a much higher accuracy at steady-state. All experimentally measured control responses, to the same reference step of $y_{\mathrm{ref}}=8$ mm, are shown opposite each other in Figure 6.
@@
From the response of the NLG-control, one can recognize a nearly uniform behavior of $x_{2}(t)$ during the whole transient phase, cf. Figure 2 (a). Note that the corresponding inclination of the $x_{1}(t)$ slope depends on the assigned $\gamma$ parameter. It is also worth noting that a residual non-zero, however minor, steady-state error of the NLG-control is due to the sensor bias and noise, while the latter is additionally impacting the used $x_{2}(t)$ signal, which is obtained via differentiation and filtering of the measured $x_{1}(t)$ value. The experimentally obtained signal $x_{2}(t)$ (when using the NLG controller) is also shown in Figure 7 for the sake of clarity.
Figure 7: The velocity state $x_{2}(t)$ of the NLG control experiment.
@@ V Summary
A novel robust nonlinear state feedback control for perturbed second-order systems was introduced. The control law combines the ideas of the state-dependent nonlinear damping introduced in [15] and the sliding mode control theory, which equips the overall feedback law with an output-sign-dependent term able to compensate for bounded perturbations. The resulted control has only one tuning parameter, the overall output feedback gain, whose selection should be based on the upper bound of the perturbation magnitude, and this closed analytic dependence (even though conservative) was shown in the paper. We provided a detailed analysis of uniform boundedness and finite-time convergence of the control system based on the proposed strict Lyapunov function. The control signal appears to be continuous during the transients, apart from the convergence instant, where the state value reaches the origin (similarly to some high order sliding mode control algorithms). Remarkable features of the proposed control are that: (i) it is equivalent (from the structural viewpoint) to a standard linear PD feedback control and, at the same time, (ii) it can guarantee (without additional integral action) for full compensation of the properly bounded matched disturbances in the second-order systems without overshoot, and arbitrary bounded in the steady state, while (iii) providing a global finite-time convergence. The proposed control was already evaluated experimentally on an electro-magneto-mechanical actuator system with considerable bounded disturbances. Analysis of its robustness abilities against process and measurement noise and development of the methods for its discretization can be considered as directions of future research.
Acknowledgments
This work was supported by AURORA mobility program (RCN grant number 340782).